OpenStudy (anonymous):

8 bit = 1 byte 1024 bytes = 1 kilobyte 1024 kilobyte = 1 megabyte 1024 megabyte = 1 gigabyte after a bit it goes with a 1024 multiplication... bit>byte>kilobyte>megabyte>gigabyte>terabyte>petabyte how many gigabytes is in 1 petabyte?

OpenStudy (anonymous):

that's wrong

OpenStudy (anonymous):

it's bit>byte>kilobyte>megabyte>gigabyte>tetrabyte>petabyte

OpenStudy (anonymous):

@jovel, no its completely right, I'm a web developer i know 6+ computer languages

OpenStudy (anonymous):

its tera not tetra, https://answers.yahoo.com/question/index?qid=20120522055559AAYJzr3

OpenStudy (johnweldon1993):

We know that 1 petabyte would be 10^15 or 1024^5 And 1 gigabyte would be 10^13 or 1024^3... so divide them \[\large \frac{1024^5}{1024^3} = 1024^2 = \space ?\]

OpenStudy (anonymous):

ok i have learnt that its bit>byte>KB>MB>GB>Terabyte>Petabyte but since you know 6+ languages i dunno

OpenStudy (eli_moses):

computer languages

OpenStudy (anonymous):

-_- oops forgot kilobyte

OpenStudy (anonymous):

corrected

OpenStudy (anonymous):

@johnweldon1993 is right

OpenStudy (anonymous):

thanks